# YSO Jets Driven by Magnetic Pressure Generated through Stellar Magnetosphere–Disk Interaction

### v1

@@ 4.1 Load: charging $B_{\phi}$ by star-disk differential rotation
On the scale shown in Figure 6, the jet is driven by magnetic pressure force primarily along the polar cavity-disk boundary, with a strong toroidal magnetic field. Since a toroidal field can in principle be generated wherever there is differential rotation along a field line, the most preferred field lines for toroidal field generation are those threading both the star, where the magnetically-dominated segment of the field line rooted on the star surface co-rotates with the star, and the disk, where the gas dominates and the field lines are forced to move with the gas at the local Keplerian velocity. We will first focus on these field lines, followed by a generalization to all field lines that experience field-gas differential rotation.
The field lines threading both the star and the dense disk generate a toroidal field in situ by a change in the angular speed of the rotation along the field line, as shown in Figure 7(d), which presents the ratio between the local angular velocity $\Omega$ and the Keplerian angular velocity $\Omega_{K}$. Across the $\beta_{K}=1$ boundary, the angular velocity transitions from nearly Keplerian in the disk to super-Keplerian above the disk. It is indicative of the differential rotation along the field line that leads to the toroidal field generation through field line twisting.
[…]
@@
To showcase a magnetic field participating in this process, Figure 7 shows a 2D projection of a 3D field line, highlighted in solid magenta. One end of this field line is anchored on the star, which transfers angular momentum to the gas via a strong magnetic field in the low plasma-$\beta$ polar region to force the gas to rotate at a super-Keplerian rate. The other end is embedded in the gas-dominated disk, where the gas motion remains nearly Keplerian. The significant differential rotation along this field line, particularly near where the field bends sharply, generates an azimuthal magnetic field, as shown in Figure 7(e).
The differential rotation along this field line generates the azimuthal magnetic field required for jet launching by magnetic pressure, accelerating the gas in the vertical $\hat{z}$-direction. Consequently, this azimuthal magnetic field is constantly advected away by gas motion. Quantitatively, magnetic field generation and removal in this nearly ideal MHD region (due to ionization by high temperature) are governed by the induction equation (equ. 3). In a cylindrical coordinate, the change in azimuthal magnetic field ($\big{[}\frac{\partial B_{\phi}}{\partial t}\big{]}_{\mathrm{tot}}$) can be divided into two processes: generation of $B_{\phi}$ from $B_{z}$ and $B_{R}$ by differential rotation,
@@
where a positive value denotes an increase in the absolute value of the toroidal magnetic field strength and a negative value denotes a decrease. Figure 7(f), (g), and (h) shows $\big{[}\frac{\partial|B_{\phi}|}{\partial t}\big{]}_{\mathrm{tot}}$, $\big{[}\frac{\partial|B_{\phi}|}{\partial t}\big{]}_{\mathrm{gen}}$, and $\big{[}\frac{\partial|B_{\phi}|}{\partial t}\big{]}_{\mathrm{adv}}$ respectively. Although $|B_{\phi}|$ decreases in magnitude in a large part of the jet-launching region at the time shown, $|B_{\phi}|$ switches rapidly between net gain and loss of strength over time, resulting in a quasi-steady state. The reader is encouraged to view the animated version of Figure 7 to see this rapid change between gaining and losing $|B_{\phi}|$.
The rapid change in the net change in $\big{[}\frac{\partial|B_{\phi}|}{\partial t}\big{]}_{\mathrm{tot}}$ does not translate to a rapid change in the $\big{[}\frac{\partial|B_{\phi}|}{\partial t}\big{]}_{\mathrm{gen}}$ nor $\big{[}\frac{\partial|B_{\phi}|}{\partial t}\big{]}_{\mathrm{adv}}$. Generation through differential rotation ($\big{[}\frac{\partial|B_{\phi}|}{\partial t}\big{]}_{\mathrm{gen}}$) stays mostly positive on the magnetically dominated side of the $\beta_{K}=1$ boundary, and advection through gas motion ($\big{[}\frac{\partial|B_{\phi}|}{\partial t}\big{]}_{\mathrm{adv}}$) stays mostly negative, illustrating the required toroidal magnetic field for jet-launching is generated in situ by differential rotation along magnetic field lines.
The $\big{[}\frac{\partial|B_{\phi}|}{\partial t}\big{]}_{\mathrm{gen}}$ term being predominantly positive in the jet-launching region indicates that toroidal field generation is not limited to field lines like the magenta fieldline highlighted in Figure 7, where one footpoint is anchored in the disk and the other on the star. In fact, any differential rotation between the dense disk gas and the stellar magnetosphere can generate a toroidal magnetic field capable of launching an outflow, even if the field line is not fully embedded in the disk. An example of this is shown in Figure 8, where a segment along an otherwise open field line becomes kinetically dominated ($\beta_{K}>1$) as it approaches the disk. This local interaction between the disk and the stellar field generates differential rotation, which in turn produces the toroidal field responsible for jet launching.
@@
Before discussing the release of toroidal magnetic energy during the “fire” phase, we first highlight that the most rapid generation of toroidal field occurs in the polar region near the star. Although this may seem counterintuitive, since field lines in the magnetically dominated polar region are typically expected to remain relatively straight, this rapid winding arises because the star rotates faster along these field lines than the disk does. Specifically, for the “two-legged” field lines illustrated in Figures 7 and 8, the footpoint on the disk lies beyond the corotation radius of the star (approximately 0.1 au). As a result, the disk end of the field line rotates more slowly than the stellar end, causing the star to drag the field line forward. This differential rotation leads to significant field twisting near the stellar surface. Although the magnetic deformation is quickly removed by outgoing Alfvén waves (and their associated Poynting flux; see Figure 9 [e, f]) and does not remain near the star, it contributes to the magnetic energy buildup that eventually drives the jet, as we describe in the next subsection.
4.2 Fire: releasing magnetic energy by global field geometry
@@
For a magnetic force to support jet launching in both hemispheres, the condition $\frac{z}{|z|}J_{R}B_{\phi}>0$ must be satisfied, such that the Lorentz force points away from the midplane. This is generally satisfied when $J_{R}>0$, given the sign of $B_{\phi}$ in Figure 7[e], which is positive in the polar cavity of the upper hemisphere and negative in that of the lower hemisphere. While $J_{R,2}$ contributes positively to $J_{R}$ and aligns with the outflow direction, $J_{R,1}$ tends to oppose the jet by introducing a restoring force toward the midplane. Moreover, $|J_{R,1}|$ is typically smaller than $|J_{R,2}|$, so the net current remains positive in the jet-launching region, as shown by the azimuthally-averaged $J_{R}$ in Figure 9(c).
This jet-driving current is generated locally along the surface of the closed magnetosphere and the disk as a consequence of the rapid changes in the toroidal magnetic field geometry in the vertical direction. As discussed in Sec. 4.1, differential rotation between the star and the disk twists magnetic field lines most effectively when the field line has one end rooted in the star and the other embedded in the disk. A comparatively strong toroidal field is produced in this narrow layer where such “two-legged” field lines reside. Above this region, in the low-density polar cavity, the strong magnetic field keeps the fieldlines roughly poloidal, and the toroidal component remains weak. Below this region is the mass-dominated disk, where the magnetic effect is more diminished, providing a solid foundation for the outflow.
As a result, the field lines within this intermediate layer, subject to significant differential rotation, are continually over-twisted relative to those above and below. This localized over-twisting establishes steep vertical gradients in the toroidal magnetic field, giving rise to a strong radial current sheet ($J_{R}$) essential for jet launching. Because this differential twisting is inherently highly variable in space and time, we illustrate it in Figure 9(d) using the azimuthally and temporally averaged quantity $\left[\frac{\partial B_{\phi}}{\partial t}\right]_{\mathrm{gen}}$. The colormap highlights the toroidal magnetic field’s generation rate in the layer above the closed magnetosphere and disk atmosphere, with more positive values indicating faster local toroidal field generation. The peaks in the toroidal field generation rate, pointed by arrows in Figure 9(d), align with the location of the $J_{R}$ current sheet at the disk surface. However, as shown in Figure 9(d), the gas in the most rapid toroidal field generating region near the star is infalling, rather than outflowing (the outflow region is outlined in white in Figure 9[d]). This raises a natural question: how does the energy generated in this inner region reach the outflow zone to drive the jet?
@@
[…]
4.3 Reload: replenishing jet gas by magnetic reconnection
The “two-legged” magnetic field lines, such as those shown in Figure 7(e), stretch and may open during outflow launching. To sustain a continuous jet (Figure 1), these field lines must reconnect and revert to a “two-legged” configuration, “reloading” with mass by magnetic reconnection between oppositely directed field lines. As discussed in Sections 3.1 and 3.3, the global magnetic topology at late times when a persistent outflow is launched consists of downward-pointing original magnetospheric field lines that are opened but remain connected to the star (referred to as “the opened stellar field lines” for short) and upward-pointing original magnetospheric field lines that are opened but thread through the (inner) disk (but not the star). These oppositely directed open field lines are expected to preferentially reconnect at their interface. Although individual reconnection events are difficult to isolate in a highly variable 3D system, we present qualitative evidence below to confirm reconnection along this boundary.
@@ 4.3 Reload: replenishing jet gas by magnetic reconnection
where constant $\Phi_{B,r}$ contours approximate field line locations, as indicated by the black dashed curves in Figure 11. These contours represent azimuthally averaged field configurations, capturing only the global magnetic evolution. In reality, three-dimensional reconnection events occur on much shorter timescales. Between $t=3.7~{}\mathrm{yr}$ (panel [a]) and $t=4.0~{}\mathrm{yr}$ (panel [d]), the same averaged field line transitions from an extended state to a more compact (lower altitude) configuration. Given the continuous outflow during this interval, the contraction of the averaged field line suggests magnetic reconnection.
Reconnection primarily occurs along the $B_{r}=0$ boundary (highlighted by the magenta contour in Figure 11) where the open field lines threading the star and disk meet. This process regulates the magnetic configuration at high altitudes, maintaining a quasi-steady state in the opened stellar magnetic flux (Figure 3[a]). In three dimensions, magnetic field stretching and localized reconnection events occur simultaneously at different heights and across multiple azimuthal sectors, ensuring quasi-continuous jet ejection.
@@
To summarize, jet launching operates through a “load-fire-reload” cycle driven by differential rotation between the star and disk. The twisting of magnetospheric field lines amplifies the toroidal magnetic field, generating a strong current sheet that accelerates the jet via the magnetic pressure force. Reconnection near the field reversal interface then resets the system, reforming field lines connecting the star and disk and reloading them with disk material for the next cycle. This process sustains a continuous outflow.
5 Discussion
5.1 Magnetospheric accretion and outflow
@@
Figure 12(a) shows a 3D rendering of the accretion streams onto the star (the red sphere at the center) The accretion streams are filamentary and follow the magnetospheric field lines. These “finger-like” accretion patterns result from the “interchange instability,” a form of Rayleigh-Taylor instability in a magnetically dominated environment. Gas from the mass-dominated disk is loaded onto the strong magnetic field and subsequently accreted onto the star following the stellar magnetic field lines. This mechanism is established by theory (J. Arons & S. M. Lea, 1976; H. C. Spruit et al., 1995) and simulations (M. M. Romanova et al., 2008; A. A. Blinova et al., 2016; S. Takasao et al., 2022; Z. Zhu et al., 2024).
@@
First, we follow Z. Zhu (2025) and define the corotation radius analytically using the stellar parameters. The corotation radius is given by
@@
where $B_{\star}$ and $R_{\star}$ are the magnetic field strength on the stellar surface and radius of the star, respectively, $M_{\star},\dot{M}$ are the stellar mass and mass accretion rate, respectively. $\bar{m}$ is the coefficient used in setting the strength of the stellar magnetosphere (equ. 8). Note that our expression is slightly different from the one in Z. Zhu (2025) because of the difference in the unit system. Using our simulation parameters and $\dot{M}=10^{-7}M_{\odot}~{}\mathrm{yr}^{-1}$ for the accretion rate at late times ($t\gtrsim 3$ yr) when the system settles to a quasi-steady state (see below, and Figure 12), the truncation radius $R_{T,\mathrm{ana}}\approx 0.088~{}\mathrm{au}$.
@@
Second, we follow S. Takasao et al. (2022) and define $R_{T}$ as the smallest radius where $\beta_{K}=1$. In our model, $R_{T,~{}\beta_{K}}$ settles to a value $\approx 0.082~{}\mathrm{au}(\approx 5R_{\star})$ at late times, slightly lower than the analytic estimate. Since the accretion stream is filamentary (Figure 12[a]), it is not surprising that the material with $\beta_{K}=1$ can penetrate slightly into the magnetosphere. The truncation radius is comparable to the corotation radius $\approx 0.1~{}\mathrm{au}$. Our $R_{T}$ is larger than S. Takasao et al. (2022) compared to the stellar radius, since our stellar magnetic field is stronger than theirs.
@@
Most of the mass is accreted along magnetic field lines onto the star away from the midplane. To illustrate the landing locations of the gas on the star, we show in Figure 12(b) the averaged mass accretion rate per unit area at different polar angles $\frac{\partial\dot{M}}{\partial A}(\theta)$ between 3.0 yr and 5.5 yr in our model measured right outside $r_{\mathrm{fix}}$ (sec. 2) at $0.02~{}\mathrm{au}$. Most accretion occurs around $\sim 20^{\circ}$-$40^{\circ}$ from the polar axis, while the accretion around the midplane ($\theta=90^{\circ}$) is negligible.
@@
The highly filamentary high-altitude accretion does not imply a strong variation in the accretion rate. Figure 12(c) shows that throughout the later stage of the simulation between 3.0 yr and 5.5 yr, the total mass accretion rate remains around $1\times 10^{-7}~{}\mathrm{M_{\odot}~{}yr^{-1}}$. In comparison, the jet outflow rate, defined as the total mass flux at ±$\pm 5~{}\mathrm{au}$ with $v_{z}^{p}>10^{7}~{}\mathrm{cm~{}s^{-1}}$, is about $2\times 10^{-9}M_{\odot}~{}\mathrm{yr^{-1}}$, around $2\%\sim 5\%$ of the gas accretion rate. The total mass outflow rate, measured at ±$\pm 5~{}\mathrm{au}$ with gas velocity $>10^{6}~{}\mathrm{cm~{}s^{-1}}$, is around $4\times 10^{-8}M_{\odot}~{}\mathrm{yr^{-1}}$. The total outflow rate is about $40\%$ of the total accretion rate, similar to the ratio in S. Takasao et al. (2022) and Z. Zhu et al. (2024), although the exact values differ mainly due to the difference in the chosen stellar magnetosphere magnetic field strength.
@@
The outflow carries angular momentum. To connect the angular momentum transported with the disk Kerperian angular momentum, we define the “Keplerian angular momentum radius” following Y. Tu et al. (2025a)
@@
where $\Phi_{l}$ and $\Phi_{m}$ are the angular momentum flux and mass flux in the outflow, respectively. Figure 12(d) shows $R_{a}$ for the jet and the disk wind, respectively, measured at $z=5$ au. The jet $R_{a}\approx 0.6~{}\mathrm{au}$, much larger than its launch radius $\lesssim 0.2~{}\mathrm{au}$. This shows an abundance of angular momentum carried by the jet, initially stored in the electromagnetic form associated with the twisting of the field lines. The disk wind $R_{a}\approx 1.3~{}\mathrm{au}$, slightly larger than its launching radius $\lesssim 1~{}\mathrm{au}$.
@@
Both the jet and the disk wind carry significant angular momentum. In the following section (sec. 5.2), we will compare the accretion and outflow properties with those driven by the disk through the avalanche-accretion stream. Since the key distinction between the disk-driven outflow and our model is the presence of a stellar magnetosphere, we will discuss its impact on the overall dynamics of the system.

### v2

@@ IV.1 Load: charging $B_{\phi}$ by star-disk differential rotation
On the scale shown in Figure 6, the jet is driven by toroidal magnetic pressure primarily along the polar cavity–disk boundary. These toroidal fields arise where differential rotation occurs along a field line, most effectively on lines threading both the star, where the magnetically dominated segment co-rotates with the stellar surface, and the disk, where the gas forces the field to move at the local Keplerian velocity (the “two-legged” field lines). These field lines connecting the star and dense disk generate toroidal fields through angular velocity gradients along their length, as illustrated in Figure 7(d). Across the $\beta_{K}=1$ boundary, the angular velocity shifts from nearly Keplerian in the disk to super-Keplerian above it, highlighting the shear that twists field lines and generates a toroidal magnetic field.
To showcase a “two-legged” magnetic field participating in this process, Figure 7 presents a 2D projection of a representative field line (solid magenta). Anchored on the star, the line transfers angular momentum to polar gas in the low-$\beta$ region, forcing super-Keplerian rotation. Its other end remains embedded in the nearly Keplerian disk. The resulting differential rotation rapidly produces a strong azimuthal field [Fig. 7(e)].
Figure 7: Toroidal magnetic field generation in the jet-launching region. The quantities plotted are azimuthally averaged. Panel (a) and (b) show the density and projected $\hat{z}-$direction velocity ($v_{z}^{p}$, equ. 9) respectively, showing the density structure and the location of the jet. Panel (c) shows the kinetic-$\beta$ (equ. 13), highlighting the transition from the magnetically-dominated polar cavity to the mass-dominated disk. Panel (d) shows the angular velocity, normalized by the local Keplerian value. The lower panels show toroidal magnetic field dynamics. Panel (e) shows the toroidal magnetic field; Panels (f), (g), and (h) show the total change in toroidal field strength (equ. 22), toroidal field generation by differential rotation (equ. 20), and toroidal field advection by flux-freezing (equ. 21) respectively. The black dashed contour highlights the $\beta_{K}=1$ surface. The magenta and green field lines are 2D projections of representative 3D field lines shown in Figure 8. An animated version is available at https://figshare.com/s/e90cdd7f3a49d07ffbc1. The movie is 25 seconds long, showing the gas properties on the upper panels, and the toroidal field dynamics on the lower panels.
@@
The strong azimuthal magnetic field empowers a jet, accelerating the gas in the vertical $\hat{z}-$direction. Consequently, this azimuthal magnetic field is constantly advected away by gas motion. Quantitatively, magnetic field generation and removal in this nearly ideal MHD region (due to ionization by high temperature) are governed by the induction equation (equ. 3). In a cylindrical coordinate, the change in azimuthal magnetic field ($\big[\frac{\partial B_{\phi}}{\partial t}\big]_{\mathrm{tot}}$) can be divided into two processes: generation of $B_{\phi}$ from $B_{z}$ and $B_{R}$ by differential rotation,
@@
where a positive value denotes an increase in the absolute value of the toroidal magnetic field strength and a negative value denotes a decrease. Figure 7(f), (g), and (h) shows $\big[\frac{\partial|B_{\phi}|}{\partial t}\big]_{\mathrm{tot}}$, $\big[\frac{\partial|B_{\phi}|}{\partial t}\big]_{\mathrm{gen}}$, and $\big[\frac{\partial|B_{\phi}|}{\partial t}\big]_{\mathrm{adv}}$ respectively [^4] Note that the flow quantities shown in Fig. 7 are first computed in each cell and then averaged azimuthally. Non-azimuthally averaged quantities yield qualitatively similar results but are much noisier.. Although $|B_{\phi}|$ decreases in magnitude in a large part of the jet-launching region at the time shown, $|B_{\phi}|$ switches rapidly between net gain and loss of strength over time, resulting in a quasi-steady state. The reader is encouraged to view the animated version of Figure 7 to see this rapid change between gaining and losing $|B_{\phi}|$.
The rapid change in the net change in $\big[\frac{\partial|B_{\phi}|}{\partial t}\big]_{\mathrm{tot}}$ does not translate to a rapid change in the $\big[\frac{\partial|B_{\phi}|}{\partial t}\big]_{\mathrm{gen}}$ nor $\big[\frac{\partial|B_{\phi}|}{\partial t}\big]_{\mathrm{adv}}$. Generation through differential rotation ($\big[\frac{\partial|B_{\phi}|}{\partial t}\big]_{\mathrm{gen}}$) stays mostly positive on the magnetically dominated side of the $\beta_{K}=1$ boundary, and advection through gas motion ($\big[\frac{\partial|B_{\phi}|}{\partial t}\big]_{\mathrm{adv}}$) stays mostly negative, illustrating the required toroidal magnetic field for jet-launching is generated in situ by differential rotation along magnetic field lines. Quantitatively, the magnitude of $\big[\tfrac{\partial|B_{\phi}|}{\partial t}\big]_{\mathrm{gen}}$ reaches $\sim 10^{3}~\mathrm{G~yr^{-1}}$ along the jet-launching disk surface, corresponding to a toroidal field generation time scale $10^{-2}~\mathrm{year}$. In comparison, the stellar surface poloidal dipole magnetic field strength is of the order of $10^{3}~\mathrm{G}$, suggesting that the toroidal field generation and advection proceed on remarkably short timescales.
The $\big[\frac{\partial|B_{\phi}|}{\partial t}\big]_{\mathrm{gen}}$ term being predominantly positive in the jet-launching region indicates that toroidal field generation is not limited to field lines like the magenta fieldline highlighted in Figure 7, where one footpoint is anchored in the disk and the other on the star. In fact, any differential rotation between the dense disk gas and the stellar magnetosphere can generate a toroidal magnetic field capable of launching an outflow, even if the field line is not fully embedded in the disk. An example of this “touch-and-go” two-legged field lines is shown in Figure 8, where a segment along an otherwise open field line becomes kinetically dominated ($\beta_{K}>1$) as it approaches the disk. This “touch-and-go” interaction is facilitated by the elevated disk atmosphere, which allows the field line to interact with the disk surface without the need to reach close to the disk midplane. The local interaction between the disk and the stellar field generates differential rotation, which in turn produces the toroidal field responsible for jet launching.
@@
IV.2 Fire: releasing magnetic energy by global field geometry
@@
For a magnetic force to support jet launching in both hemispheres, the condition $\frac{z}{|z|}J_{R}B_{\phi}>0$ must be satisfied, such that the Lorentz force points away from the midplane. This is generally satisfied when $J_{R}>0$, given the sign of $B_{\phi}$ in Figure 7[e], which is positive in the polar cavity of the upper hemisphere and negative in that of the lower hemisphere in jet-launching inner disk region. While $J_{R,1}$ tends to oppose the jet by introducing a restoring force toward the midplane, $J_{R,2}$ contributes positively to $J_{R}$, overcoming the restoring force by $J_{R,1}$, and aligning with the outflow direction.
This jet-driving current is generated locally along the surface of the closed magnetosphere and the disk as a consequence of the rapid changes in the toroidal magnetic field geometry in the vertical direction. As discussed in Sec. IV.1, the toroidal field is most rapidly generated when the field line has one end rooted in the star and the other embedded in the disk. A comparatively strong toroidal field is produced in this narrow layer where such “two-legged” field lines reside. Above this region, in the low-density polar cavity, the strong magnetic field keeps the fieldlines roughly poloidal, and the toroidal component remains weak. Below this region is the mass-dominated disk, where the magnetic effect is more diminished, providing a solid foundation for the outflow.
As a result, the field lines within this intermediate layer, subject to significant differential rotation, are continually over-twisted relative to those above and below. This localized over-twisting establishes steep vertical gradients in the toroidal magnetic field, giving rise to a strong radial current sheet ($J_{R}$) essential for jet launching. Because this differential twisting is inherently highly variable in space and time, we illustrate it in Figure 9(d) using the azimuthally and temporally averaged quantity $\left[\frac{\partial B_{\phi}}{\partial t}\right]_{\mathrm{gen}}$. The colormap highlights the extremely rapid generation rate of the toroidal magnetic field ($\gtrsim 10^{3}~\mathrm{G~yr^{-1}}$) in the layer above the closed magnetosphere and disk atmosphere, with more positive values indicating faster local toroidal field generation. The peaks in the toroidal field generation rate, pointed by arrows in Figure 9(d), align with the location of the $J_{R}$ current sheet at the disk surface. However, as shown in Figure 9(d), the gas in the most rapid toroidal field generating region near the star is infalling, rather than outflowing (the outflow region is outlined in white in Figure 9[d]). This raises a natural question: how does the energy generated in this inner region reach the outflow zone to drive the jet?
@@
Figure 10: 3D rendition of the protostar, accreting magnetosphere, and magnetic field lines. The disk is shown as semi-transparent density isosurfaces at around $10^{-12}~\mathrm{g~cm^{-3}}$. The protostar is represented by the central red solid sphere. White lines depict magnetic field lines, including permanently closed magnetospheric loops and twisted field lines resulting from magnetosphere–disk interactions (e.g., the prominent twisted fieldline near the top-center of the image). An animated version of this figure is available at: https://figshare.com/s/413012328cf225f86a67. The movie is 27 seconds long, showing the magnetic field lines near the magnetosphere.
IV.3 Reload: replenishing jet gas by magnetic reconnection
The “two-legged” magnetic field lines, such as those shown in Figure 7(e), stretch and may open during outflow launching. To sustain a continuous jet (Figure 1), these field lines must reconnect and revert to a “two-legged” configuration, “reloading” with mass by magnetic reconnection between oppositely directed field lines. As discussed in Sections III.1 and III.3, the global magnetic topology at late times when a persistent outflow is launched consists of downward-pointing original magnetospheric field lines that are opened but remain connected to the star (referred to as “the opened stellar field lines” for short) and upward-pointing original magnetospheric field lines that are opened but thread through the (inner) disk (but not the star). These oppositely directed open field lines are expected to preferentially reconnect at their interface. Although individual reconnection events are difficult to isolate in a highly variable and rapidly evolving 3D system, we present qualitative evidence below to confirm reconnection along this boundary.
@@ IV.3 Reload: replenishing jet gas by magnetic reconnection
where constant $\Phi_{B,r}$ contours approximate field line locations, as indicated by the black dashed curves in Figure 11. These contours represent azimuthally averaged field configurations, capturing only the global magnetic evolution. In reality, three-dimensional reconnection events occur on much shorter timescales. Between $t=3.7~\mathrm{yr}$ (panel [a]) and $t=4.0~\mathrm{yr}$ (panel [d]), the same averaged field line transitions from an extended state to a more compact (lower altitude) configuration. Given the continuous outflow during this interval, the contraction of the averaged field line suggests magnetic reconnection, which primarily occurs along the $B_{r}=0$ boundary (highlighted by the magenta contour in Figure 11) where the open field lines threading the star and disk meet. Crucially, a full reset of the magnetic geometry into a closed dipolar loop is not required to sustain the jet: the required toroidal field can be generated by differential rotation as long as a star-anchored fieldline can become partially and shallowly embedded in the gas-dominated, magnetically elevated disk, as shown in sec. IV.2 and Fig. 8. Consequently, individual reconnection events occur more rapidly and on much shorter time scales than in 2D models, making the so-called “magnetospheric ejections (MEs)” associated with prominent coherent reconnection events in 2D (e.g., Zanni and Ferreira, 2013) harder to identify in 3D. How these 2D MEs are modified by severe tangling of magnetic field lines in the turbulent (MRI-active) and magnetically puffed-up inner disk remains to be quantified.
@@
To summarize, jet launching operates through a “load-fire-reload” cycle driven by differential rotation between the star and disk. The twisting of magnetospheric field lines amplifies the toroidal magnetic field, generating a strong current sheet that accelerates the jet via the magnetic pressure force. Reconnection near the field reversal interface then resets the system, reforming field lines connecting the star and disk and reloading them with disk material for the next cycle. In our 3D simulation, individual reconnection events occur asynchronously across azimuth—that is, concurrently but without global coherence, with different $\phi$ sectors loading and releasing mass and magnetic energy independently. This spatial and temporal asynchrony sustains a rather continuous jet despite the intrinsically episodic nature of each local event.
V Discussion
V.1 Magnetospheric accretion and outflow
@@
Figure 10 shows a 3D rendering of the accretion streams onto the star (the red sphere at the center) The accretion streams are filamentary and follow the magnetospheric field lines. These “finger-like” accretion patterns result from the “interchange instability,” a form of Rayleigh-Taylor instability in a magnetically dominated environment. Gas from the mass-dominated disk is loaded onto the strong magnetic field and subsequently accreted onto the star following the stellar magnetic field lines. This mechanism is established by theory (Arons and Lea, 1976; Spruit et al., 1995) and simulations (Romanova et al., 2008; Blinova et al., 2016; Takasao et al., 2022; Zhu et al., 2024).
@@
First, we follow Zhu (2025) and define the truncation radius analytically using the stellar and disk accretion parameters [^5] We note that there is a long history of analytically estimating the magnetospheric truncation radius, as discussed in, e.g., Bessolaz et al. (2008). The analytic expression we adopted from Zhu (2025) is identical to that recommended by Bessolaz et al. (2008) (see their equation [6]).:
@@
where $B_{\star}$ and $R_{\star}$ are the magnetic field strength on the stellar surface and radius of the star, respectively, $M_{\star},\dot{M}$ are the stellar mass and mass accretion rate, respectively. $\bar{m}$ is the coefficient used in setting the strength of the stellar magnetosphere (equ. 8). Using our simulation parameters and $\dot{M}=10^{-7}M_{\odot}~\mathrm{yr}^{-1}$ for the accretion rate at late times ($t\gtrsim 3$ yr) when the system settles to a quasi-steady state (see below, and Figure 12), the truncation radius $R_{T,\mathrm{ana}}\approx 0.088~\mathrm{au}$.
@@
Second, we follow Takasao et al. (2022) and define $R_{T}$ as the smallest radius where $\beta_{K}=1$. In our model, $R_{T,~\beta_{K}}$ settles to a value $\approx 0.082~\mathrm{au}(\approx 5R_{\star})$ at late times, slightly lower than the analytical estimate. Since the accretion stream is filamentary (Figure 10), it is not surprising that the material with $\beta_{K}=1$ can penetrate slightly into the magnetosphere. The truncation radius is comparable to the corotation radius $\approx 0.1~\mathrm{au}$. Our $R_{T}$ is larger than Takasao et al. (2022) compared to the stellar radius, since our stellar magnetic field is stronger than theirs. Using either definition, the truncation radius is smaller than the corotation radius, which is slightly larger than 0.1 au.
@@
Most of the mass is accreted along magnetic field lines onto the star away from the midplane. To illustrate the landing locations of the gas on the star, we show in Figure 12(a) the averaged mass accretion rate per unit area at different polar angles $\frac{\partial\dot{M}}{\partial A}(\theta)$ between 3.0 yr and 5.5 yr in our model measured right outside $r_{\mathrm{fix}}$ (sec. II) at $0.02~\mathrm{au}$. Most accretion occurs around $\sim 20^{\circ}$-$40^{\circ}$ from the polar axis, while the accretion around the midplane ($\theta=90^{\circ}$) is negligible.
@@
The highly filamentary high-altitude accretion does not imply a strong variation in the accretion rate. Figure 12(b) shows that throughout the later stage of the simulation between 3.0 yr and 5.5 yr, the total mass accretion rate remains around $1\times 10^{-7}~\mathrm{M_{\odot}~yr^{-1}}$. In comparison, the jet outflow rate, defined as the total mass flux at $\pm 5~\mathrm{au}$ with $v_{z}^{p}>10^{7}~\mathrm{cm~s^{-1}}$, is about $2\times 10^{-9}M_{\odot}~\mathrm{yr^{-1}}$, around $2\%\sim 5\%$ of the gas accretion rate. The total mass outflow rate, measured at $\pm 5~\mathrm{au}$ with gas velocity $10^{6}~\mathrm{cm~s^{-1}}$, is around $4\times 10^{-8}M_{\odot}~\mathrm{yr^{-1}}$. The total outflow rate is about $40\%$ of the total accretion rate, similar to the ratio in Takasao et al. (2022) and Zhu et al. (2024), although the exact values differ mainly due to the difference in the chosen stellar magnetosphere magnetic field strength.
@@
The outflow carries an angular momentum. To connect the angular momentum transported by the outflow with the disk Keplerian angular momentum, we define the “Keplerian angular momentum radius” following Tu et al. (2025a)
@@
where $\Phi_{l}=\int\rho lv_{z}dA$ ($l$ is the specific angular momentum) and $\Phi_{m}=\int\rho v_{z}dA$ are the angular momentum flux and mass flux in the outflow, respectively. Figure 12(c) shows $R_{a}$ for the jet and the disk wind, respectively, measured at $z=5$ au. The jet $R_{a}\approx 0.6~\mathrm{au}$, much larger than its launch radius $\lesssim 0.2~\mathrm{au}$. This shows an abundance of angular momentum carried by the jet, initially stored in the electromagnetic form associated with the twisting of the field lines. The disk wind $R_{a}\approx 1.3~\mathrm{au}$, slightly larger than its launching radius $\lesssim 1~\mathrm{au}$.
@@
Both the jet and the disk wind carry significant angular momentum. In the following section (sec. V.2), we will compare the accretion and outflow properties with those driven by the disk through the avalanche-accretion stream. Since the key distinction between disk-driven outflow and our model is the presence of a stellar magnetosphere, we will discuss its impact on the overall dynamics of the system.